OpenStudy (anonymous):

Five of Ethan's family members went on a 4 and one-sixth hour road trip. If they all shared the drive time equally, how many hours did each family member drive?

OpenStudy (mathstudent55):

\( 4 \frac{1}{6} \div 5 \)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[4\frac{ 1 }{ 6 }\div5=\frac{ 5 }{6 } or .8333\]

OpenStudy (anonymous):

Thanks nathan

OpenStudy (mathstudent55):

\(4 \dfrac{1}{6} \div 5 \) \(=4 \dfrac{1}{6} \times \dfrac{1}{5} \) \( = \dfrac{25}{6} \times \dfrac{1}{5} \) Can you finish it now?

OpenStudy (anonymous):

yeah

OpenStudy (mathstudent55):

Great.